How to prove the identity (tanu-1)/(tanu+1)= (1-cotu)/(1+cotu)
Answer by lwsshak3(11628)

How to prove the identity
(tanu-1)/(tanu+1)= (1-cotu)/(1+cotu)
start with left side
(tan-1)/(tan+1)
=(1/cot-1)/(1/cot+1)
=(1-cot/cot)/(1+cot/cot)
=1-cot/1+cot
verified: left side=right side